OpenStudy (anonymous):

4x-3/6 + 2>=2x-1/12. solve and write in interval

OpenStudy (anonymous):

\[4x-1/2+2\ge2x-1/12\] \[4x-2x+3/2\ge-1/12\] \[2x \ge-1/12-3/2\] \[2x \ge-19/12\] \[x \ge-19/24\]

OpenStudy (anonymous):

Thanks idk why I keep messing it up